OpenStudy (anonymous):

integral of xsqrt(x^2 + 1)

OpenStudy (zepp):

\[\large \int x\sqrt{x^2+1}dx\]Could you try to apply some substituion here? :D

OpenStudy (anonymous):

I'm really bad at u substitutions. I always get confused as to which is substituted. Is it u=x^2+1 and du=2x? then I replace the x^2+1 with u and x with du? But then do I need to put a 1/2 out infront of the integral?

OpenStudy (anonymous):

use \[u=x^2+1\] \[du=2x\] You will need to multiply and divide by 2 and you're set.

OpenStudy (anonymous):

is the final answer 1/3(x^2+1)^(3/2) +c?????? :)

OpenStudy (zepp):

Basically, the substitution rule is to take a complicated thing and use a letter to define that, and then, you could use derivative to get the dx out and substitution wrt the letter you assigned to that thing

OpenStudy (anonymous):

Yes it is Heather, that's right.

OpenStudy (anonymous):

Yay! Thank you both so much!

OpenStudy (zepp):

\[\large \int x\sqrt{x^2+1}dx \\\text{Let} ~u=x^2+1;~~du = 2x~dx \\\frac{du}{2}=xdx \\\large \int \frac{1}{2}\sqrt{u}du=\frac{1}{2}\int \sqrt{u}du=\frac{1x^{\frac{3}{2}}}{3}+c=\frac{(x^2+1)^{\frac{3}{2}}}{3}+c \]yup, correct :)

OpenStudy (anonymous):

Thank you!!